OpenStudy (anonymous):

Use the law of cosines to solve the triangle: a=16km B=40 degrees c=10km

OpenStudy (anonymous):

Do you know the law of cosine?

OpenStudy (anonymous):

Just learned it today

OpenStudy (anonymous):

One second I'm going to work it out

OpenStudy (jdoe0001):

\(\large \bf \textit{Law of Cosines}\\ \quad \\ c^2 = a^2+b^2-(2ab)cos(c)\\ c = \sqrt{a^2+b^2-(2ab)cos(c)}\)

OpenStudy (anonymous):

Ok

OpenStudy (jdoe0001):

do you have a picture? and what solve, means? find side b?

OpenStudy (anonymous):

I have to find the missing sides and angles. How do u insert a picture

OpenStudy (jdoe0001):

[Attache File] blue buttton, to the left-side of [Post] blue button

OpenStudy (anonymous):

I don't see it

OpenStudy (anonymous):

\[b^2=16^2+10^2-2(10)(16)*\cos(40)\]Square the 16 and the 10. Also multiply the -2 with (10)(16) Which gives you \[b^2=356-320*\cos(40)\] Do you know how to solve for b from here?

OpenStudy (anonymous):

Yes

OpenStudy (jdoe0001):

@Laurene96 well, you can post it at say http://imgur.com and give us the url to it

OpenStudy (anonymous):

I got 5.25

OpenStudy (anonymous):

Ok I'm uploading the pic

OpenStudy (anonymous):

http://imgur.com/gbOpQ5j It's number 6

OpenStudy (anonymous):

Type the whole "356-320*cos(40)" into your calculator. You subtracted first then did the cos

OpenStudy (anonymous):

Oh ok. One sec.

OpenStudy (anonymous):

10.53

OpenStudy (anonymous):

That's right! I'm guessing they want you to round to the nearest tenth? so 10.5

OpenStudy (anonymous):

She usually wants us to do hundredths but I'm not sure

OpenStudy (jdoe0001):

\(\bf \textit{Law of Cosines}\\ \quad \\ b^2 = {\color{blue}{ a}}^2+{\color{green}{ c}}^2-(2{\color{blue}{ a}}{\color{green}{ c}})cos(B)\\ b = \sqrt{{\color{blue}{ a}}^2+{\color{green}{ c}}^2-(2{\color{blue}{ a}}{\color{green}{ c}})cos(B)}\\ \quad \\ b = \sqrt{{\color{blue}{ 16}}^2+{\color{green}{ 10}}^2-(2\cdot {\color{blue}{ 16}}\cdot {\color{green}{ 10}})cos(40^o)}\)

OpenStudy (jdoe0001):

anyhow.. as Ryan88 has shown :)

OpenStudy (anonymous):

Well it's probably best to do as your teacher likes, so b = 10.53

OpenStudy (jdoe0001):

you have to find the angles A and C, right?

OpenStudy (anonymous):

Alright

OpenStudy (anonymous):

And yes

OpenStudy (anonymous):

Would I use sin-1 to find angle A?

OpenStudy (jdoe0001):

well, you'd use the Law of Cosines as well... lemme .. try instead say angle C \(\bf \textit{Law of Cosines}\\ \quad \\ c^2 = a^2+b^2-(2ab)cos(C)\implies c^2-a^2-b^2=-(2ab)cos(C)\\ \quad \\ a^2+b^2-c^2=(2ab)cos(C)\implies \cfrac{a^2+b^2-c^2}{(2ab)}=cos(C)\\ \quad \\ {\color{red}{ \square }}=cos(C)\implies cos^{-1}({\color{red}{ \square }})=cos^{-1}[cos(C)]\implies cos^{-1}({\color{red}{ \square }})=\measuredangle C\)

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Use the Law of Sin \[\frac{ 10.53 }{ \sin(40) }=\frac{ 16 }{ Sin(A) }\]

OpenStudy (anonymous):

Either or, both do the exact same

OpenStudy (anonymous):

I got 77.61

OpenStudy (jdoe0001):

for the angle?

OpenStudy (anonymous):

Yeah

OpenStudy (jdoe0001):

\(\bf \cfrac{16^2+10.53^2-10^2}{2(16)(10.53)}=cos(C)\\ \quad \\\implies 37.62480093287110170194 \approx cos(C)\)

OpenStudy (jdoe0001):

wel.... one...hmm ..lemme recheck that

OpenStudy (jdoe0001):

\(\bf \cfrac{16^2+10.53^2-10^2}{2(16)(10.53)}=cos(C)\\ \quad \\\implies 0.79202546296296296296\approx cos(C)\\ \quad \\37.62480093287110170194 \approx \measuredangle C\)

OpenStudy (anonymous):

Oh I was looking for angle a instead of c

OpenStudy (anonymous):

I got 37.62 for angle C.

OpenStudy (anonymous):

And 102.38 for angle A

OpenStudy (jdoe0001):

I got 37.62 for angle C And 102.38 for angle A \(\Large \checkmark\)

OpenStudy (anonymous):

Thanks for the help!

OpenStudy (jdoe0001):

yw